OpenStudy (anonymous):

solve for x. 5/(2x+3)=4/x can someone show me the process because there is a lot of these ones

HanAkoSolo (jamierox4ev3r):

\(\frac{5}{2x+3}=\frac{4}{x}\)

HanAkoSolo (jamierox4ev3r):

if i understand correctly, is what the problm looks like @5occ3r ?

OpenStudy (anonymous):

first you would solve the problem is to isolate the variable x

OpenStudy (anonymous):

yes @Jamierox4ev3r

OpenStudy (anonymous):

first cancel out fractions so how would you go about isolating this | v 5 / (2x+3) = 4 / x

HanAkoSolo (jamierox4ev3r):

5x=8x+12 -8x -8x ------- -3x=12 /-3 /-3 ------- x=-4

OpenStudy (anonymous):

>.<

HanAkoSolo (jamierox4ev3r):

just cross multiply the two fraction to create one simple fraction is all you have to do

HanAkoSolo (jamierox4ev3r):

sorry @mth3v4 but i was solving first >.<

OpenStudy (anonymous):

lol i think my way is obsolete by now

HanAkoSolo (jamierox4ev3r):

so does that make sense @5occ3r ? btw nice way of spelling soccer :D

OpenStudy (anonymous):

how did you get 5x=8x+12? and thanks :D

OpenStudy (anonymous):

@Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

by cross multiplying. 5*x=5x, and 4*2x+3=8x+12 Therefore, 5x=8x+12. Does that make sense?

OpenStudy (anonymous):

Yes thank you! @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

no problem im glad i could explain it to you :D